# Submitted toTopology Proceedings

### v2

@@ 3. Global elliptic sectors
Let $x\in\partial\mathcal{S}\setminus\{a\}$ and suppose that $\Gamma(x)\subset\partial\mathcal{S}\setminus\{a\}$ is bounded. The set $K:=\overline{\Gamma(x)}$ is compact and $\Gamma(x)\cup w_{\pm}(\Gamma(x))\subset K\subset\partial\mathcal{S}$. By [10, Theorem 5.1], $\Gamma(x)$ either is a periodic orbit with exactly one equilibrium, a center $\tilde{a}$, in $\operatorname{Int}(\Gamma(x))$, or $w_{+}(\Gamma(x))$ and $w_{-}(\Gamma(x))$ each consist of exactly one equilibrium. If the first case occurs, by openness of the basin $\mathcal{V}$ in $\tilde{a}$, cf. Theorem 2.4, there would be a point $y\in\mathcal{S}\cap\mathcal{V}$, $y\not=x$. Thus, $\Gamma(y)$ would be periodic with $a$ in its limit sets. This is impossible and the second case in [10, Theorem 5.1] must occur.
By Proposition 3.5, $\Gamma(x)$ must be a homoclinic orbit with $w_{\pm}(\Gamma(x))=\{a\}$. In particular, by [10, Proposition 4.3 (ii)], $\Gamma(x)$ tends to $a$ in definite directions $\eta_{\pm}\in\mathcal{E}(F,m)$ for $t\to\pm\infty$. Moreover, there exist $r,\Theta_{\pm}>0$ such that for all $z_{\pm}\in\{z\in\mathbb{C}{}:|z-a|<r,|\arg(z-a)-\eta_{\pm}|<\Theta_{\pm}\}$ the orbit through $z_{\pm}$ tends to $a$ in the definite direction $\eta_{\pm}$ for $t\to\pm\infty$. Additionally, there exists $\tau>0$ such that $M_{\pm}:=\Phi(\pm\tau,x)\in A_{\pm}$. Choose $\varepsilon_{\pm}>0$ small enough such that $\mathcal{B}_{\varepsilon_{\pm}}(M_{\pm})\subset A_{\pm}$. Since $M_{\pm}\in\partial\mathcal{S}$, there exists $y_{\pm}\in\mathcal{B}_{\varepsilon_{\pm}}(M_{\pm})\cap\mathcal{S}$. As an orbit cannot tend to $a$ in more than one direction for $t\to\infty$ and $t\to-\infty$, it follows $\theta_{+}=\eta_{+}$ and $\theta_{-}=\eta_{-}$, i.e. $\Gamma(x)$ must be $(\theta_{+},\theta_{-})$-elliptic in $a$.
Define $\varepsilon:=\min\{\varepsilon_{+},\varepsilon_{-}\}$. Let $L_{0}$ be the curve segment of $\Gamma(x)$ from $M_{-}$ to $M_{+}$. As in Step 2 of the proof of Theorem 3.6, we choose paths without contact $l_{\pm}\subset\mathcal{B}_{\varepsilon}(M_{\pm})$ through $M_{\pm}$ and construct an elementary topological quadrangle $R$ around $L_{0}$ defined via the functions $\chi_{1}$, $\chi_{2}$ and $\zeta$. Define $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}(\theta_{+},\theta_{-},\Xi)$. Since $x\in\partial\mathcal{S}$ is an interior point of $R$ and $\operatorname{Ext}(\Xi\cup\{a\})$, there exists $\hat{y}\in\mathcal{S}^{\prime}\cap R$, $\Gamma(\hat{y})\not=\Gamma(x)$.
If $\hat{y}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$, we would have $\Xi\subset\operatorname{Int}(\Gamma(\hat{y})\cup\{a\})\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and thus $x\in\partial\mathcal{S}\cap\mathcal{S}$, which is a contradiction to Theorem 3.6. Hence we must have $\hat{y}\in\operatorname{Ext}(\Gamma(x)\cup\{a\})$. Choose $\Gamma_{1}=\Gamma(x)$ and $\Gamma_{2}=\Gamma(\hat{y})\subset\operatorname{Ext}(\Gamma_{1}\cup\{a\})$ in (IMEO) to conclude $\Gamma(x)\subset\operatorname{Int}(\Gamma(\hat{y})\cup\{a\})$. We get $\Xi\subset\operatorname{Ext}(\Gamma(x)\cup\{a\})\cap\operatorname{Int}(\Gamma(\hat{y})\cup\{a\})$. Additionally, both orbits, $\Gamma(x)$ and $\Gamma(\hat{y})$ cross $l_{-}\subset\partial R$. As in Step 4 of the proof of Theorem 3.6, it follows $\Xi\cap R\not=\emptyset$. By choosing $\Gamma_{1}=\Gamma(x)$ and $\Gamma_{2}=\Xi$ in (IMEO), we conclude again $\Xi\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$ and thus $x\in\partial\mathcal{S}\cap\mathcal{S}$. This is again a contradiction to Theorem 3.6. Thus, $\Gamma(x)$ cannot be homoclinic and must be unbounded. It follows that also $\mathcal{S}$ is unbounded. ∎
@@
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. It holds
@@
Let $x\in\mathcal{S}$. Suppose there exists $z_{1}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$ such that $z_{1}\not\in\mathcal{S}$. Then there exists $z_{2}\in\partial\mathcal{S}\cap\operatorname{Int}(\Gamma(x)\cup\{a\})\not=\emptyset$. Since the case $a=z_{2}$ is not possible, $\Gamma(z_{2})$ has to be unbounded, cf. Theorem 3.7. But $\Gamma(z_{2})$ is path-connected and thus must lie completely in $\operatorname{Int}(\Gamma(x)\cup\{a\})$, which is bounded. This is a contradiction and $z_{1}$ cannot exist. Hence we have $\operatorname{Int}(\Gamma(x)\cup\{a\})\subset\mathcal{S}$. Additionally, by applying Proposition 3.4, we get
These considerations show both equations in (3.1).
Suppose there exist $y_{1},y_{2}\in\mathcal{S}$, $\Gamma(y_{1})\not=\Gamma(y_{2})$, such that $\Gamma(y_{1})$ and $\Gamma(y_{2})$ are mutually exterior, i.e. $\Gamma(y_{1})\subset\operatorname{Ext}(\Gamma(y_{2})\cup\{a\})$ and $\Gamma(y_{2})\subset\operatorname{Ext}(\Gamma(y_{1})\cup\{a\})$.
At first, assume that $\{y_{1},y_{2}\}\cap\operatorname{Ext}(\Xi\cup\{a\})\not=\emptyset$. To fix ideas, assume w.l.o.g. $y_{1}\in\operatorname{Ext}(\Xi\cup\{a\})$. We have $\Xi\subset\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$. Since $\Gamma(y_{2})\subset\operatorname{Ext}(\Gamma(y_{1})\cup\{a\})$, we get $y_{2}\in\operatorname{Ext}(\Xi\cup\{a\})$ and thus $\Xi\subset\operatorname{Int}(\Gamma(y_{1})\cup\{a\})\subset\operatorname{Ext}(\Gamma(y_{2})\cup\{a\})$. It follows the contradiction $y_{2}\not\in\mathcal{S}$.
Secondly, we assume $\{y_{1},y_{2}\}\subset\operatorname{Int}(\Xi\cup\{a\})$. By using (3.1), we have $\operatorname{Int}(\Gamma(y_{1})\cup\{a\})\subset\mathcal{S}$, i.e. all orbits in $\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$ are $(\theta_{+},\theta_{-})$-elliptic in $a$ and $\hat{\Xi}:=\Gamma(y_{1})$ is another sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$. Define $\hat{\mathcal{S}}:=\mathcal{S}(\theta_{+},\theta_{-},\hat{\Xi})$. Since $\Xi$ is $(\theta_{+},\theta_{-})$-elliptic and $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})$, we get $\Xi\subset\hat{\mathcal{S}}$. By using (3.1), we conclude $y_{2}\in\operatorname{Int}(\Xi\cup\{a\})\subset\hat{\mathcal{S}}$. Additionally, since $y_{2}\in\operatorname{Ext}(\hat{\Xi}\cup\{a\})$, we must have $\Gamma(y_{1})=\hat{\Xi}\subset\operatorname{Int}(\Gamma(y_{2})\cup\{a\})$, which is a contradiction.
The last possible case is $\{y_{1},y_{2}\}\cap\Xi\not=\emptyset$. To fix ideas, we assume w.l.o.g. $\Xi=\Gamma(y_{1})$. By Definition 3.1 (iii), we must have either $y_{2}\in\operatorname{Int}(\Xi\cup\{a\})=\operatorname{Int}(\Gamma(y_{1})\cup\{a\})$ or $\Gamma(y_{1})=\Xi\subset\operatorname{Int}(\Gamma(y_{2})\cup\{a\})$. Both results lead to a contradiction.
Finally, $y_{1}$ and $y_{2}$ cannot exist and we conclude that all orbits in $\mathcal{S}$ must be nested. ∎
@@
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. $\mathcal{S}$ is connected, path-connected and simply connected.
@@
Suppose, there exists a separation $(U,V)$ of $\mathcal{S}$, i.e. $\mathcal{S}\subset U\cup V$, where $U$ and $V$ are open, disjoint and have nonempty intersection with $\mathcal{S}$. Then $a$ can only be in one of these two sets. Assume w.l.o.g. $a\in U$. Choose $y\in V\cap\mathcal{S}$. Since $\Gamma(y)$ is connected, we get $\Gamma(y)\subset V$. Since $a\not\in\overline{V}$, it follows the contradiction $a\not\in w_{+}(\Gamma(y))\cup w_{-}(\Gamma(y))$. Hence, such a separation cannot exist and $\mathcal{S}$ is connected. By [23, Proposition 12.25], $\mathcal{V}$ is even path-connected.
Let $L\subset\mathcal{S}$ be a loop. Define $J_{x}:=\Gamma(x)\cup\{a\}$ for $x\in\mathcal{S}$. Due to the geometry described in Proposition 3.8, we can use the same idea as in the proof of Theorem 2.5. By Theorem 3.6, for all $x\in L$ exists $y_{x}\in\operatorname{Ext}(J_{x})\cap\mathcal{S}$ such that $x\in\operatorname{Int}(J_{y_{x}})$. We can replace the periodic orbits $\Gamma(y_{x})$, $x\in\mathcal{V}$, by the curves $J_{y_{x}}$, $x\in\mathcal{S}$. Thus, we can find a suitable finite open cover of $L$ and an outermost orbit $J\subset\mathcal{S}$ satisfying $a\in J$ and $L\subset\operatorname{Int}(J)$. By (3.1), we have $\operatorname{Int}(J\cup\{a\})\subset\mathcal{S}$, which is clearly simply connected. Hence $L$ can be continuously deformed to a point in $L\subset\mathcal{S}$. Since $\mathcal{S}$ is path-connected, it follows that $\mathcal{S}$ is even simply connected. ∎
We now address the independence of the global elliptic sector from a particular choice of the sector-forming orbit.
@@
Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. For all $x\in\mathcal{S}$ the orbit $\Gamma(x)$ is a sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ and $\mathcal{S}=\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))$, i.e. the global elliptic sector $\mathcal{S}$ does not depend on the particular choice of a sector-forming orbit.
@@
Let $x\in\mathcal{S}$ be arbitrary. By (3.1), all orbits in $\operatorname{Int}(\Gamma(x)\cup\{a\})$ are $(\theta_{+},\theta_{-})$-elliptic and $\hat{\Xi}:=\Gamma(x)$ is indeed an sector-forming orbit in $a$ with definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$.
Define $\hat{\mathcal{S}}:=\mathcal{S}(\theta_{+},\theta_{-},\hat{\Xi})$. If $\hat{\Xi}=\Xi$, then obviously $\mathcal{S}=\hat{\mathcal{S}}$. We have to show $\mathcal{S}\subset\hat{\mathcal{S}}$ as well as $\hat{\mathcal{S}}\subset\mathcal{S}$ for the two cases $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})$ and $\Xi\subset\operatorname{Int}(\hat{\Xi}\cup\{a\})$. As these four proofs are all based on similar arguments, we assume w.l.o.g. $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})$ and show only $\mathcal{S}\subset\hat{\mathcal{S}}$. Let $y\in\mathcal{S}$. If $y\in\Xi\cup\hat{\Xi}\cup\operatorname{Int}(\hat{\Xi}\cup\{a\})$, then nothing is to show. Thus, assume $y\in\operatorname{Ext}(\hat{\Xi}\cup\{a\})\setminus\Xi$. If $y\in\operatorname{Ext}(\Xi\cup\{a\})$, then we have $\Xi\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$ and thus $\hat{\Xi}\subset\operatorname{Int}(\Xi\cup\{a\})\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$, i.e. $y\in\hat{\mathcal{S}}$. Additionally, if $y\in\operatorname{Ext}(\hat{\Xi}\cup\{a\})\cap\operatorname{Int}(\Xi\cup\{a\})$, then we can apply Proposition 3.8 with $y_{1}:=x$ and $y_{2}:=y$ to conclude $\hat{\Xi}\subset\operatorname{Int}(\Gamma(y)\cup\{a\})$, i.e. $y\in\hat{\mathcal{S}}$. ∎
Finally, at the end of this chapter, we want to supplement and enhance our topological study with two important corollaries, which confirm our intuitive idea of the geometrical structure of an equilibrium of order $m\geq 2$, cf. [10, Figure 2, Figure 3].
@@
There exist exactly $22m-$ 2 distinct444in the sense of [2, §17, 3.]. global elliptic sectors in $a$. For all $x\in\mathbb{C}{}$ with $w_{+}(\Gamma(x))=w_{-}(\Gamma(x))=\{a\}$ there exist two adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $\Gamma(x)$ is a sector-forming orbit for $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))$.
@@
By [10, Theorem 4.4], there is a FED of order $d:=2m-2$ in $a$. We use the geometrical objects $\Gamma$, $\Gamma_{i}$ and $p_{i}$, $i\in\{1,\ldots,d\}$, defined by [10, Definition 4.1 d)]. We order the definite directions $\mathcal{E}(F,m)=\{\theta_{1},\ldots,\theta_{d}\}$ in such a way that for all $i\in\{1,\ldots,d\}$ the local elliptic sector with characteristic orbits $\Gamma_{i}$ and $\Gamma_{i+1}$ lies between $\theta_{i}$ and $\theta_{i+1}$, i.e. $\Gamma_{i}$ tends to $a$ in the definite direction $\theta_{i}$. [^5] If $i=d$, we define $\Gamma_{i+1}:=\Gamma_{1}$ and $\theta_{i+1}:=\theta_{1}$. By using the notation in [10, Definition 4.1 a)] and applying [10, Definition 4.1 c) (iv)], there exist points $E_{i}\in\Gamma(p_{i},p_{i+1})$ such that $\Xi_{i}:=\Gamma(E_{i})$ is a sector-forming orbit with definite directions $\theta_{i},\theta_{i+1}$ for all $i\in\{1,\ldots,d\}$. Moreover, since all local sectors in $a$ have pairwise empty intersection except for the characteristic orbits and $a$, these sector-forming orbits are pairwise mutually exterior. Define the set
$A:=\overline{\operatorname{Int}(\Gamma)}\setminus\bigcup_{i=1}^{d}\overline{\operatorname{Int}(\Xi_{i}\cup\{a\})}$
as well as the $d$ global elliptic sectors
$\mathcal{S}_{i}:=\begin{dcases}\mathcal{S}(\theta_{i},\theta_{i+1},\Xi_{i})&\text{if }\lambda(\theta_{i})=-1\\ \mathcal{S}(\theta_{i+1},\theta_{i},\Xi_{i})&\text{if }\lambda(\theta_{i})=1\end{dcases}\;,\,i\in\{1,\ldots,d\}\text{.}$
Suppose there exist $i,j\in\{1,\ldots,d\}$, $i\not=j$, with $z\in\mathcal{S}_{i}\cap\mathcal{S}_{j}\not=\emptyset$. Since $\Xi_{i}$ and $\Xi_{j}$ are mutually exterior, we must have $z\in\operatorname{Ext}(\Xi_{i}\cup\{a\})\cap\operatorname{Ext}(\Xi_{j}\cup\{a\})$ and thus $\Xi_{i}\cup\Xi_{j}\subset\operatorname{Int}(\Gamma(z)\cup\{a\})$. By (3.1), we have $\operatorname{Int}(\Gamma(z)\cup\{a\})\subset\mathcal{S}_{i}$, i.e. $\Xi_{i}\cup\Xi_{j}\subset\mathcal{S}_{i}$. Since $\Xi_{i}$ and $\Xi_{j}$ are mutually exterior, we can derive a contradiction to the second statement in Proposition 3.8. Thus, these global elliptic sectors are distinct.
Let $x\in\mathbb{C}{}\setminus\{a\}$ be arbitrary with $w_{+}(\Gamma(x))=w_{-}(\Gamma(x))=\{a\}$. By [10, Proposition 4.3 (ii)], $\Gamma(x)$ has to tend to $a$ in definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ for $t\to\infty$ and $t\to-\infty$, respectively.
At first, we assume $\Gamma(x)\subset\overline{\operatorname{Int}(\Gamma)}$. Then, by applying the geometrical structure of the FED and [10, Proposition 4.3 (iii)], all orbits through $A$ leave $\overline{\operatorname{Int}(\Gamma)}$ in finite time. Hence, there exists $i\in\{1,\ldots,d\}$ such that $\Gamma(x)\subset\Xi_{i}\cup\operatorname{Int}(\Xi_{i}\cup\{a\})\subset\mathcal{S}_{i}$. By Definition 3.1 (iii), it follows that $\{\theta_{+},\theta_{-}\}=\{\theta_{i},\theta_{i+1}\}$, i.e. $\theta_{+},\theta_{-}$ are adjacent.
Next, we consider the case $\Gamma(x)\cap\operatorname{Ext}(\Gamma)\not=\emptyset$. As $\Gamma(x)$ must leave and enter $\overline{\operatorname{Int}(\Gamma)}$ again, a part of the closed Jordan curve $\Gamma$ must lie in $\operatorname{Int}(\Gamma(x)\cup\{a\})$. By [10, Definition 4.1 c) (iii)], there must exist $i\in\{1,\ldots,d\}$ with $E_{i}\in\operatorname{Int}(\Gamma(x)\cup\{a\})$, i.e. $\Xi_{i}\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$. Suppose a second sector-forming orbit is in the interior of $\Gamma(x)\cup\{a\}$, i.e. there exists $j\not=i$ with $\Xi_{j}\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$. If $x\in\mathcal{S}_{i}$, then, by applying the second statement in Proposition 3.8, $\Xi_{i}$ and $\Xi_{j}$ cannot be mutually exterior, which is a contradiction. Thus, $x\not\in\mathcal{S}_{i}$. Since $\mathcal{S}_{i}$ is connected, cf. Theorem 3.9, we conclude $\mathcal{S}_{i}\subset\operatorname{Int}(\Gamma(x)\cup\{a\})$. This makes $\mathcal{S}_{i}$ bounded, which is a contradiction to Theorem 3.7. Hence, $\Xi_{i}$ is the only sector-forming orbit in $\operatorname{Int}(\Gamma(x)\cup\{a\})$. By applying the geometrical structure of the FED, we again get $\{\theta_{+},\theta_{-}\}=\{\theta_{i},\theta_{i+1}\}$, i.e. $\theta_{+},\theta_{-}$ are adjacent. In particular, by Definition 3.1 (iii), also in this case it follows $x\in\mathcal{S}_{i}$.
Finally, in both cases we can apply Lemma 3.10 to conclude that $\Gamma(x)$ is a sector forming orbit for the global elliptic sector $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))$ and $\mathcal{S}(\theta_{+},\theta_{-},\Gamma(x))=\mathcal{S}_{i}$. This shows the second statement. Moreover, using this result, there cannot exist more then the $d$ distinct global elliptic sectors $\mathcal{S}_{1},\ldots,\mathcal{S}_{d}$. This completes the proof of the first statement. ∎
@@
We have proved Corollary 3.11 by applying the geometry of the local structure in $a$. This structure is defined by a FED, which is not uniquely determined, i.e. there exist several FEDs in $a$. Nevertheless, we can now see that the global structure is indeed unique.
@@
Let $m\geq 3$. Let $\mathcal{S}:=\mathcal{S}(\theta_{+},\theta_{-},\Xi)$ be a global elliptic sector with adjacent definite directions $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$, $\lambda(\theta_{\pm})=\mp 1$, and a sector-forming orbit $\Xi$. Then it holds
(3.2) $\mathcal{S}(\theta_{+},\theta_{-},\Xi)=\left\{x\in\mathbb{C}{}:\Gamma(x)\text{ is }(\theta_{+},\theta_{-})\text{-elliptic in }a\right\}\text{.}$

### v3

@@ 3.2. The geometry of global elliptic sectors
Let $x\in\partial\mathcal{S}\setminus\{a\}$ and suppose that $\Gamma(x)\subset\partial\mathcal{S}\setminus\{a\}$ is bounded. The set $K:=\overline{\Gamma(x)}$ is compact and $\Gamma(x)\cup w_{\pm}(\Gamma(x))\subset K\subset\partial\mathcal{S}$. By using the holomorphy of $F$, we can apply Theorem 1.1 to conclude that $\Gamma(x)$ either is a periodic orbit with exactly one equilibrium, a center $\tilde{a}$, in $\operatorname{Int}(\Gamma(x))$, or $w_{+}(\Gamma(x))$ and $w_{-}(\Gamma(x))$ each consist of exactly one equilibrium. If the first case occurs, by openness of the basin $\mathcal{V}$ in $\tilde{a}$, cf. Theorem 2.4, there would be a point $y\in\mathcal{S}\cap\mathcal{V}$, $y\not=x$. This is impossible and the second case in Theorem 1.1 must occur. Hence, by Proposition 3.4, $\Gamma(x)$ must be a homoclinic orbit with $w_{\pm}(\Gamma(x))=\{a\}$. Lemma 3.5 gives us a homoclinic orbit $\hat{\Gamma}$ in $\mathcal{S}$ such that $\Gamma(x)$ and $\hat{\Gamma}$ are nested. However, by openness of $\mathcal{S}$, this is impossible. Thus, $\Gamma(x)$ cannot be homoclinic and must be unbounded. It follows that also $\mathcal{S}$ is unbounded. ∎
@@
Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. It holds
@@
This proof requires several case distinctions and is quite technical. A proof can be found in the appendix. ∎
@@
Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. $\mathcal{S}$ is connected, path-connected and simply connected.
@@
Analogous to the proof of Theorem 2.5, this follows directly from (3.1) together with the Seifert-van Kampen theorem for infinite open covers, cf. [15, Theorem 1.20]. ∎
We now address the independence of the global elliptic sector from a particular choice of the sector-forming orbit ensuring in some sense the well-definedness of the global elliptic sector.
@@
Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. For all $x\in\mathcal{S}$ the orbit $\Gamma(x)$ is a sector-forming orbit in $a$ and $\mathcal{S}=\mathcal{S}(\Gamma(x))$, i.e. the global elliptic sector $\mathcal{S}$ does not depend on the particular choice of a sector-forming orbit.
@@
Let $x\in\mathcal{S}$. By (3.1), all orbits in $\operatorname{Int}(\Gamma(x)\cup\{a\})$ are homoclinic and $\hat{\Xi}:=\Gamma(x)$ is indeed a sector-forming orbit in $a$.
[…]
3.3. Final Corollaries
Finally, at the end of this chapter, we supplement and enhance our topological study with several important corollaries, which confirm the intuitive idea of the geometrical structure of an equilibrium of order $m\geq 2$, cf. [19, Figure 2, Figure 3].
@@ 3.3. Final Corollaries
There exist exactly 2m22m-2 distinct [^4] This means that $\Gamma_{1}$ and $\Gamma_{2}$ are mutually exterior, if if they belong to two different global elliptic sectors. global elliptic sectors in $a$, each located between two adjacent definite directions given by $\mathcal{E}(F,m)$. All homoclinic orbits are sector-forming orbits.
@@
By [19, Theorem 4.4], there is a FED of order d:=2m2d:=2m-2 in $a$. We use the geometrical objects $\Gamma$, $\Gamma_{i}$ and $p_{i}$, $i\in\{1,\ldots,d\}$, defined by [19, Definition 4.1 d)], cf Figure 1. We order the definite directions $\mathcal{E}(F,m)=\{\theta_{1},\ldots,\theta_{d}\}$ in such a way that for all $i\in\{1,\ldots,d\}$ the local elliptic sector with characteristic orbits $\Gamma_{i}$ and $\Gamma_{i+1}$ lies between $\theta_{i}$ and $\theta_{i+1}$, i.e. $\Gamma_{i}$ tends to $a$ in the definite direction $\theta_{i}$. By using [19, Definition 4.1 c) (iv)], there exist points $E_{i}\in\Gamma(p_{i},p_{i+1})$ such that $\Xi_{i}:=\Gamma(E_{i})$ is a sector-forming orbit with definite directions $\theta_{i},\theta_{i+1}$ for all $i\in\{1,\ldots,d\}$. Moreover, since all local sectors in $a$ have pairwise empty intersection except for the characteristic orbits and $a$, these sector-forming orbits are pairwise mutually exterior. This leads to $d$ global elliptic sectors, generated by the $d$ pairwise mutually exterior sector-forming orbits $\Xi_{1},\ldots,\Xi_{d}$. On the one hand, due to the local geometry of $a$ it is clear that each global elliptic sector admits at least one of the orbits $\Xi_{1},\ldots,\Xi_{d}$. On the other hand, since all orbits in a global elliptic sector are nested, it is not possible to find two orbits $\Xi_{i},\Xi_{j}$, $i\not=j$, lying in the same global elliptic sector. Hence, we get indeed exactly $d$ distinct global elliptic sectors in $a$. The rest follows from the geometry described in [19, Proposition 4.3]. ∎
@@
We have proved Corollary 3.11 by using the geometry of the local structure in $a$. This structure is defined by a FED, which is not uniquely determined, i.e. there exist several FEDs in $a$. However, we can now see that the global structure is the unique globalization of this local structure.
In the following, we consider the case $m\geq 3$. In this setting, the global elliptic sector can be defined independently of sector-forming orbits, solely by means of the chosen adjacent definite directions. For this purpose, we require the function $\lambda:\mathcal{E}(F,m)\to\{-1,1\}$,
$\lambda(\theta):=\cos(\operatorname{arg}(F^{(m)}(a))+\theta(m-1))$
to determine the time direction (positive or negative) in which the orbits tend to $a$ in the FED, cf. [19, Proposition 4.3 (iii)] as well as the calculation of $\lambda_{i}$ on page 11 in this reference.
Definition 3.12.
Let $m\geq 3$.
  • (i)
    Let $\Gamma\subset\mathbb{C}$ be an orbit and $x\in\Gamma$. We call $\Gamma$ $(\theta_{+},\theta_{-})$-elliptic in $a$, if there exist $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that $w_{+}(\Gamma)=w_{-}(\Gamma)=\{a\}$, $\arg(\Phi(t,x)-a)=\theta_{+}$ for $t\to\infty$ and $\arg(\Phi(t,x)-a)=\theta_{-}$ for $t\to$-$\infty$, that is, $\Gamma$ is homoclinic and tends to $a$ in the definite direction $\theta_{\pm}$ for $t\to\pm\infty$.
  • (ii)
    Let $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ be two adjacent definite directions satisfying $\lambda(\theta_{\pm})=\mp 1$. [^5] If the corresponding local elliptic sector between $\theta_{+}$ and $\theta_{-}$ has clockwise (counterclockwise) direction, then $\theta_{-(+)}=\theta_{+(-)}+\frac{\pi}{m-1}$. We define the set
    (3.2) $\mathcal{S}(\theta_{+},\theta_{-}):=\left\{x\in\mathbb{C}{}:\Gamma(x)\text{ is }(\theta_{+},\theta_{-})\text{-elliptic in }a\right\}\text{.}$
@@
Let $m\geq 3$.
  • (i)
    Let $\mathcal{S}:=\mathcal{S}(\Xi)$ be a global elliptic sector generated by the sector-forming orbit $\Xi$. Then there exist $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ such that
    (3.3) $\mathcal{S}(\Xi)=\mathcal{S}(\theta_{+},\theta_{-})\text{.}$
  • (ii)
    Let $\theta_{+},\theta_{-}\in\mathcal{E}(F,m)$ be two adjacent definite directions satisfying $\lambda(\theta_{\pm})=\mp 1$. Then there exists a sector forming orbit $\Xi$ such that
    (3.4) $\mathcal{S}(\theta_{+},\theta_{-})=\mathcal{S}(\Xi)\text{.}$